OpenStudy (anonymous):

what is the angle between v=(-7,8) and the positive x-axis? I need to use a formula (which I will write below) to find the cosine of the angle between the two vectors, (-7,8) and (1,0).

OpenStudy (anonymous):

\[v1\times v2= \left| v1 \right|\left| v2 \right|\cos \theta \]

OpenStudy (anonymous):

u dot v = ||u||*||v|| cos a where a is the angle between u and v (0<a<pi)

OpenStudy (anonymous):

don't use x because there is a cross product also. use a dot or the word dot when representing the dor product.

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{u \cdot v}{||u||\times ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\)

OpenStudy (anonymous):

if you solve for cos a and then take the inverse cos, you'll get a (theta in your equation).

OpenStudy (anonymous):

i just need help plugging the numbers into the equation

OpenStudy (anonymous):

||v|| = sqrt( (-7)^2 + (8)^2)) ||u|| = sqrt( (1)^2 + (0)^2))

OpenStudy (anonymous):

u dot v you should know from the previous problem